OpenStudy (anonymous):

simplify radicals: -3√24a^4b^2c^3

OpenStudy (luigi0210):

\[-3\sqrt{4*6*a^4*b^2*c^2*c}\]

OpenStudy (luigi0210):

simplify what's inside and that's 4*6 btw

OpenStudy (jhannybean):

\[\large -3\sqrt{24\cdot a^4\cdot b^2\cdot c^3}\]Expand this so you can pull our numbers in groups of 2, because when things are under a radical,you can pull out anything that is a group of 2. \[\large -3\sqrt{3 \cdot 2 \cdot 2\cdot 2\cdot a \cdot a \cdot a \cdot a \cdot b \cdot b \cdot c \cdot c \cdot c}\] pull out every pair of 2 similar numbers and multiply them with -3 \[\large -3 \cdot 2 \cdot a^2 \cdot b \cdot c \sqrt{6 \cdot c}\] simplify \[\large -6a^2bc \sqrt{6c}\]

OpenStudy (luigi0210):

@stephanievega she's the one with the explanations so watch out ;P

OpenStudy (anonymous):

how did you get 3 inside the radical ?

OpenStudy (jhannybean):

break apart 24 . 24 = 6 x 4 = (3 x 2) x (2 x 2)